# bibliography.bib

### v1

@@ 1 Introduction and Main Results
A spiral set is a set of the form $\{\sqrt{n}u_{n}\;|\;n\in\mathbb{N}\}$ such that $u_{n}\in\mathbb{S}^{1}$. A spiral set is called a Fermat spiral if $u_{n}=e^{2\pi i\alpha n}$ for some $\alpha\in\mathbb{R}$. Fermat spirals, and, in particular, the “sunflower spiral” (the Fermat spiral with constant angle difference $\varphi=(1+\sqrt{5})/2$, the golden ratio), have long been studied. In [Vogel_1979], Vogel studied a close relative of the sunflower spiral that is found in nature (specifically, with angle difference $\varphi^{2}$), and in [Marklof_2020], Marklof studied spiral sets where the angle is of the form $\alpha\sqrt{n}$.
@@
and there is a decreasing function $V\colon(0,\infty)\to(0,\infty)$, called a visibility function, such that for every $\varepsilon>0$, every $\varepsilon\times V(\varepsilon)$ rectangle intersects $X$. In [Akiyama_2020], Akiyama showed that the Fermat spirals with constant angle difference are Delone if and only if the angle difference is badly approximable, and in [Adiceam_2022] Adiceam and Tsokanos studied the visibility properties of general spiral sets, questions that are related to the open Danzer problem (see [Adiceam_2021]). In this paper we show that no Fermat spiral is a dense forest, which answers a question posed by Adiceam and Tsokanos whether the Fermat spiral with $\alpha=\varphi=(\sqrt{5}+1)/2$, the golden ratio, (also called the sunflower spiral) is a dense forest. In fact, we prove a stronger result about Chabauty limits of Fermat spirals. We do this by identifying all the Chabauty limits of some Fermat spirals.
[…]
@@
Let $X$ be a Fermat spiral with badly-approximable angle difference $\alpha$, and let $(\beta,c,\tilde{c})$ be the limit of some subsequence of
@@
for every such a triplet $(\beta,c,\tilde{c})$, rotation matrix $R$, and $0<t$, where $A_{t}=\left[\begin{smallmatrix}t\\ &1/t\end{smallmatrix}\right]$.
@@ 2 Applications of Theorem 1
Assume that $x\in\tilde{X}$, then for any $n$ denote by $x_{n}$ the point closest to $x$ in $T_{n}X$ (there must be such a point since $T_{n}X$ is closed). For every $10<\varepsilon<$ 1 and $\|x\|<1/\varepsilon$, for some $N$ and every $N\leq n$, we have that $d(X,T_{n}X)<\varepsilon$, therefore $x\in\mathcal{N}_{\varepsilon}(T_{n}X)$ and so $\|x_{n}-x\|<\varepsilon$.
In the other direction, we have a sequence $x_{n}\in T_{n}X$ that converges to $x$, and denote by $\tilde{x}_{n}\in\tilde{X}$ the point closest to $x_{n}$. For every $10<\varepsilon<$ 1 there is $N_{1}$ such that for every $N_{1}\leq n$ we have $\|x-x_{n}\|<\varepsilon$; similarly, there is $N_{2}$ such that for every $N_{2}\leq n$ we have $x_{n}\in\mathcal{N}_{\varepsilon}(\tilde{X})$, therefore $\|\tilde{x}_{n}-x_{n}\|<\varepsilon$. For $N=\max\{N_{1},N_{2}\}$ and $N\leq n$ we have
@@
There are exactly six closed subgroups of $\mathbb{R}^{2}$ (up to isomorphism): 00, $\mathbb{Z}$, $\mathbb{R}$, $\mathbb{Z}^{2}$, $\mathbb{Z}\times\mathbb{R}$, $\mathbb{R}^{2}$. In [Akiyama_2020], Akiyama showed that for a badly-approximable angle difference, the Fermat spiral is Delone, therefore the limits must be Delone as well, and the only Delone closed subgroups of $\mathbb{R}^{2}$ are the lattices. ∎
@@
Let $\alpha$ be the angle difference. If $\alpha\in\mathbb{Q}$ the points lie on a finite number of rays from the origin, and so for some $\varepsilon>0$ we can look at the half-infinite strip $[a,\infty)\times[\varepsilon/2,\varepsilon]$ and for large enough $a$ it will not intersect any point. If $\alpha\notin\mathbb{Q}$ is not badly-approximable, Akiyama showed in [Akiyama_2020] that the Fermat spiral is not relatively dense, and if $\alpha\notin\mathbb{Q}$ is badly-approximable, it follows from Corollary 5 and Proposition 6. ∎
@@ 3 Proof of Theorem 1
$x_{t+b_{j}}-x_{t}=$ $\sqrt{t+b_{j}}e^{2\pi i\alpha(t+b_{j})}-\sqrt{t}e^{2\pi i\alpha t}=$
$\underbrace{(\sqrt{t+b_{j}}-\sqrt{t})e^{2\pi i\alpha(t+b_{j})}}_{u_{t}^{1}}+\underbrace{\sqrt{t}(e^{2\pi i\alpha(t+b_{j})}-e^{2\pi i\alpha t})}_{u_{t}^{2}}.$
@@
For $i=1$ we have that the angle of $u_{n_{j}}^{1}$ is $2\pi\alpha(n_{j}+b_{j})$, while the angle of $u_{n_{j}+a_{j}}^{1}$ is $2\pi\alpha(n_{j}+a_{j}+b_{j})$, so the difference in the angles is $2\pi\alpha a_{j}$ (modulo $2\pi$) which must converge to 00 otherwise $x_{n_{j}+a_{j}}-x_{n_{j}}$ will not converge. The norms are $\sqrt{n_{j}+b_{j}}-\sqrt{n_{j}}$ and $\sqrt{n_{j}+a_{j}+b_{j}}-\sqrt{n_{j}+a_{j}}$ respectively. Without loss of generality, assume $|a_{j}|\leq|b_{j}|$ and $|\sqrt{n_{j}+b_{j}}-\sqrt{n_{j}}|\leq c$, therefore
@@
$d_{j}\coloneqq$ $\sqrt{n_{j}+b_{j}}-\sqrt{n_{j}}-(\sqrt{n_{j}+a_{j}+b_{j}}-\sqrt{n_{j}+a_{j}})\leq$
$\sqrt{n_{j}+|b_{j}|}-\sqrt{n_{j}}-\Bigl{(}\sqrt{n_{j}+2|b_{j}|}-\sqrt{n_{j}+|b_{j}|}\Bigr{)}\leq$
$\sqrt{n_{j}+c\sqrt{n_{j}}}-\sqrt{n_{j}}-\Bigl{(}\sqrt{n_{j}+2c\sqrt{n_{j}}}-\sqrt{n_{j}+c\sqrt{n_{j}}}\Bigr{)}.$
@@
$d_{j}\leq$ ${n_{j}}^{1/4}\biggl{(}\frac{c}{2}\cdot{n_{j}}^{-1/4}+O\bigl{(}{n_{j}}^{-1/2}\bigr{)}\biggr{)}-{n_{j}}^{1/4}\biggl{(}\frac{c}{2}\cdot{n_{j}}^{-1/4}+O\bigl{(}{n_{j}}^{-1/2}\bigr{)}\biggr{)}=$
$\frac{c}{2}-\frac{c}{2}+O\bigl{(}{n_{j}}^{-1/4}\bigr{)}=o(1).$
This means we have $d_{j}\to 0$, therefore $u^{1}_{n_{j}+a_{j}}-u^{1}_{n_{j}}\to 0$.
@@
$\lim_{j\to\infty}u_{n_{j}+a_{j}}^{2}-u_{n_{j}}^{2}=$ $\lim_{j\to\infty}(x_{n_{j}+a_{j}}-x_{n_{j}})(e^{2\pi i\alpha b_{j}}-1)=$
$v_{1}\lim_{j\to\infty}(e^{2\pi i\alpha b_{j}}-1)=v_{1}(1-1)=0.$
To show that the Chabauty limit is closed under taking inverses, we can look at the points $x_{n_{j}-a_{j}}$, and if $x_{n_{j}-a_{j}}-x_{n_{j}}$ converges to $v$, then from the argument above we have that $v+v_{1}=0$, therefore $-v_{1}=v$ is in the Chabauty limit. Note that we have
@@ 4 Properties of the Limit
converges to a point in $\mathbb{R}^{3}$ and we will denote the limit by $(\beta,c,\tilde{c})$. We have $\beta\geq 1$ and $|c|,|\tilde{c}|\geq c_{\alpha}>0$.
@@
The spiral set is relatively dense (see [Akiyama_2020]), so there is $0<r$ such that for every $j$ there is $a_{j}$ such that $\|x_{n_{j}}-x_{a_{j}}\|<r$, therefore, by the definition of $q_{j}$, $\|x_{n_{j}}-x_{q_{j}}\|<r$, and the sequence $x_{n_{j}-q_{j}}-x_{n_{j}}$ must have a converging subsequence, say to a vector $-v\neq 0$. Therefore, we have that a subsequence of $x_{n_{j}+q_{j}}-x_{n_{j}}$ converges to the vector $v$. Let $u\neq 0$ be some vector in the lattice, then there is a sequence $a_{j}$ such that $x_{n_{j}+a_{j}}-x_{n_{j}}\to u$, and by Proposition 8 for each $j$ we have
@@
Denote by $A$ the convex hull of ±$\pm v$ and ±$\pm\tilde{v}$. Assume that the lattice is not spanned by $v$ and $\tilde{v}$, so there is a vector $u\in\Lambda$ that is not spanned by them, and let $a_{j}$ be a sequence such that $x_{n_{j}+a_{j}}-x_{n_{j}}\to u$. Without loss of generality, we may assume that $u$ is in the basic parallelogram with the vertices at 00, $v$, $\tilde{v}$, and $v+\tilde{v}$. If $u\notin A$, then it is in the triangle with its vertices at $v$, $\tilde{v}$, and $v+\tilde{v}$, therefore $u-(v+\tilde{v})\in A$. This means that without loss of generality we may assume that $u\in A$ and we may also assume that $a_{j}>0$ because otherwise we can take $x_{n_{j}-a_{j}}-x_{n_{j}}\to-u\in A$.
@@
$B_{j}=\{x\in\mathbb{R}^{2}\;|\;\text{the angle between $x$ and $x_{n_{j}}$ is at most $2\pi|\alpha q_{j}-p_{j}|$}\}.$
@@
$\bigl{(}\tilde{q}_{j}/q_{j},q_{j}(q_{j}\alpha-p_{j}),\tilde{q}_{j}(\tilde{q}_{j}\alpha-\tilde{p}_{j})\bigr{)}\to(\beta,c,\tilde{c}).$
@@
$\lim_{j\to\infty}v_{j}^{2}=\lim_{j\to\infty}e^{2\pi i\alpha n_{j}}\sqrt{n_{j}+q_{j}}(e^{2\pi i\alpha q_{j}}-1)=e^{i\theta}\cdot\lim_{j\to\infty}\sqrt{n_{j}+q_{j}}q(e^{2\pi i\alpha q_{j}}-1).$
@@
$\mkern-36.0mu\operatorname{Re}\sqrt{n_{j}+q_{j}}(e^{2\pi i\alpha q_{j}}-1)=\sqrt{n_{j}+q_{j}}\cos(2\pi\alpha q_{j})=$
$\sqrt{n_{j}+q_{j}}O\biggl{(}\frac{1}{{q_{j}}^{2}}\biggr{)}=\sqrt{\frac{n_{j}}{{q_{j}^{2}}}+\frac{1}{q_{j}}}O\biggl{(}\frac{1}{q_{j}}\biggr{)}=$
$\biggl{(}\frac{n_{j}}{q_{j}}+o(1)\biggr{)}O\biggl{(}\frac{1}{q_{j}}\biggr{)}=(2t+o(1))o(1)=o(1)$
$\mkern-36.0mu\operatorname{Im}\sqrt{n_{j}+q_{j}}(e^{2\pi i\alpha q_{j}}-1)=\sqrt{n_{j}+q_{j}}\sin(2\pi\alpha q_{j})=$
$\sqrt{n_{j}+q_{j}}\biggl{(}2\pi(\alpha q_{j}-p_{j})+O\biggl{(}\frac{1}{{q_{j}}^{3}}\biggr{)}\biggr{)}=$
$\sqrt{\frac{n_{j}}{{q_{j}}^{2}}+\frac{1}{q_{j}}}\biggl{(}2\pi c+o(1)+O\biggl{(}\frac{1}{{q_{j}}^{2}}\biggr{)}\biggr{)}=$
$\biggl{(}\frac{\sqrt{n_{j}}}{q_{j}}+o(1)\biggr{)}\cdot(2\pi c+o(1))=\frac{1}{2t}\cdot 2\pi c+o(1)=\frac{\pi c}{t}+o(1),$
@@
The co-volume of the lattice $v\mathbb{Z}+\tilde{v}\mathbb{Z}$ is the volume of the fundamental parallelogram of the lattice, that is $|\det(v,\tilde{v})|$.
@@
Proposition 12 clearly follows from the main result is given in [Yamagishi_2021], where they prove that the area of Voronoi cells converges to $\pi$ (see the paper for definitions). However, we give a brief proof for completeness.
@@
$\frac{\tilde{c}}{\beta}-c\beta=$ $\bigl{(}q_{j}(q_{j}\alpha-p_{j})+o(1)\bigr{)}\biggl{(}\frac{q_{j}}{\tilde{q}}+o(1)\biggr{)}-$
$\qquad\bigl{(}\tilde{q}_{j}(\tilde{q}_{j}\alpha-\tilde{p}_{j})+o(1)\bigr{)}\biggl{(}\frac{\tilde{q}}{q_{j}}+o(1)\biggr{)}=$
$\tilde{q}_{j}(q_{j}\alpha-p_{j})-q_{j}(\tilde{q}_{j}\alpha-\tilde{p}_{j})+o(1),$
@@
$\biggl{|}\frac{\tilde{c}}{\beta}-c\beta\biggr{|}=$ $|\tilde{q}_{j}(q_{j}\alpha-p_{j})-q_{j}(\tilde{q}_{j}\alpha-\tilde{p}_{j})|+o(1)=$
$\tilde{q}_{j}|q_{j}\alpha-p_{j}|+q_{j}|\tilde{q}_{j}\alpha-\tilde{p}_{j})|+o(1)=1+o(1),$
@@
$\bigl{(}\tilde{q}_{j}/q_{j},q_{j}(q_{j}\alpha-p_{j}),\tilde{q}_{j}(\tilde{q}_{j}\alpha-\tilde{p}_{j})\bigr{)}\to(\beta,c,\tilde{c}).$
@@
$\mkern-36.0mu\operatorname{Re}e^{-2\pi i\alpha n_{j}}(x_{n_{j}+q_{j}}-x_{n_{j}})=\operatorname{Re}\sqrt{n_{j}+q_{j}}e^{2\pi i\alpha q_{j}}-\sqrt{n_{j}}=$
$\sqrt{n_{j}+q_{j}}\biggl{(}1-O\biggl{(}\frac{1}{{q_{j}}^{2}}\biggr{)}\biggr{)}-\sqrt{n_{j}}=$
$\sqrt{n_{j}+q_{j}}-\sqrt{n_{j}}+\sqrt{\frac{n_{j}}{{q_{j}}^{2}}+\frac{1}{q_{j}}}\cdot o(1)=$
$t+o(1)+\biggl{(}\frac{1}{2t}+o(1)\biggr{)}\cdot o(1)=t+o(1)$
$\mkern-36.0mu\operatorname{Im}e^{-2\pi i\alpha n_{j}}(x_{n_{j}+q_{j}}-x_{n_{j}})=\operatorname{Im}\sqrt{n_{j}+q_{j}}e^{2\pi i\alpha q_{j}}-\sqrt{n_{j}}=$
$\sqrt{n_{j}+q_{j}}\sin(2\pi\alpha q_{j})=\sqrt{n_{j}+q_{j}}\biggl{(}2\pi(q_{j}\alpha-p_{j})+O\biggl{(}\frac{1}{{q_{j}}^{3}}\biggr{)}\biggr{)}=$
$\sqrt{\frac{n_{j}}{{q_{j}}^{2}}+\frac{1}{q_{j}}}\biggl{(}2\pi c+O\biggl{(}\frac{1}{{q_{j}}^{2}}\biggr{)}\biggr{)}=\biggl{(}\frac{1}{2t}+o(1)\biggr{)}(2\pi c+o(1))=\frac{\pi c}{t}+o(1),$
@@
The author conjectures that there is a spiral set that is a dense forest, this is in contrast to a conjecture of Adiceam and Tsokanos in [Adiceam_2022]. A possible avenue to consider is noting that if all of the Chabauty limits of a set are dense forests, the set is also a dense forest (by Proposition 6), and then constructing a spiral set whose limits are similar to the construction given by Bishop in [Bishop_2011] using the lattice limits given by Fermat spirals. If there are indeed spiral sets that are dense forests, one might ask what is the best visibility function, and in particular whether there is a Danzer spiral set.

### v2

# Chabauty Limits of Fermat Spirals

@@ 1 Introduction and Main Results
A spiral set is a set of the form $\{\sqrt{n}u_{n}\;|\;n\in\mathbb{N}\}$ such that $u_{n}\in\mathbb{S}^{1}$. A spiral set is called a Fermat spiral if $u_{n}=e^{2\pi i\alpha n}$ for some $\alpha\in\mathbb{R}$. Fermat spirals, and, in particular, the sunflower spiral (the Fermat spiral with constant angle difference $\varphi=(1+\sqrt{5})/2$, the golden ratio), have long been studied. In [Vog79], Vogel studied a close relative of the sunflower spiral that is found in nature (specifically, with angle difference $\varphi^{2}$), and in [Mar20], Marklof studied spiral sets where the angle is of the form $\alpha\sqrt{n}$.
@@
and there is a decreasing function $V\colon(0,\infty)\to(0,\infty)$, called a visibility function, such that for every $\varepsilon>0$, every $\varepsilon\times V(\varepsilon)$ rectangle intersects $X$. In [Aki20], Akiyama showed that the Fermat spirals with constant angle difference are Delone if and only if the angle difference is badly approximable, and in [AT22] Adiceam and Tsokanos studied the visibility properties of general spiral sets, questions that are related to the open Danzer problem (see [Adi21]). In this paper we show that no Fermat spiral is a dense forest, which answers a question posed by Adiceam and Tsokanos whether the sunflower spiral is a dense forest. In fact, we prove a stronger result about Chabauty limits of Fermat spirals. We do this by identifying all the Chabauty limits of some Fermat spirals.
Note that in the case that $X$ is a spiral set the $\limsup$ in eq. 1 may be replaced by $\lim$, and the limit is always $1$.
@@
[…]
@@
[…]
@@ 2 Applications of Theorem 1
Assume that $x\in\tilde{X}$, then for any $n$ denote by $x_{n}$ the point closest to $x$ in $T_{n}X$ (there must be such a point since $T_{n}X$ is closed). For every $0<\varepsilon<1$ and $\|x\|<1/\varepsilon$, for some $N$ and every $N\leq n$, we have that $d(X,T_{n}X)<\varepsilon$, therefore $x\in\mathcal{N}_{\varepsilon}(T_{n}X)$ and so $\|x_{n}-x\|<\varepsilon$.
[…]
@@
There are exactly six closed subgroups of $\mathbb{R}^{2}$ (up to isomorphism): 0, $\mathbb{Z}$, $\mathbb{R}$, $\mathbb{Z}^{2}$, $\mathbb{Z}\times\mathbb{R}$, $\mathbb{R}^{2}$. In [Aki20], Akiyama showed that for a badly-approximable angle difference, the Fermat spiral is Delone, therefore the limits must be Delone as well, and the only Delone closed subgroups of $\mathbb{R}^{2}$ are the lattices. ∎
@@
Let $\alpha$ be the angle difference. If $\alpha\in\mathbb{Q}$ the points lie on a finite number of rays from the origin, and so for some $\varepsilon>0$ we can look at the half-infinite strip $a,\infty)\times[\varepsilon/2,\varepsilon]$ and for large enough $a$ it will not intersect any point. If $\alpha\notin\mathbb{Q}$ is not badly-approximable, Akiyama showed in [Aki20] that the Fermat spiral is not relatively dense, and if $\alpha\notin\mathbb{Q}$ is badly-approximable, it follows from Corollary 5 and Proposition 6. ∎
@@ 3 Proof of Theorem 1
$x_{t+b_{j}}-x_{t}$ $=\sqrt{t+b_{j}}e^{2\pi i\alpha(t+b_{j})}-\sqrt{t}e^{2\pi i\alpha t}$
$=\underbrace{(\sqrt{t+b_{j}}-\sqrt{t})e^{2\pi i\alpha(t+b_{j})}}_{u_{t}^{1}}+\underbrace{\sqrt{t}(e^{2\pi i\alpha(t+b_{j})}-e^{2\pi i\alpha t})}_{u_{t}^{2}}.$
@@
For $i=1$ we have that the angle of $u_{n_{j}}^{1}$ is $2\pi\alpha(n_{j}+b_{j})$, while the angle of $u_{n_{j}+a_{j}}^{1}$ is $2\pi\alpha(n_{j}+a_{j}+b_{j})$, so the difference in the angles is $2\pi\alpha a_{j}$ (modulo $2\pi$) which must converge to 0 otherwise $x_{n_{j}+a_{j}}-x_{n_{j}}$ will not converge. The norms are $\sqrt{n_{j}+b_{j}}-\sqrt{n_{j}}$ and $\sqrt{n_{j}+a_{j}+b_{j}}-\sqrt{n_{j}+a_{j}}$ respectively. Without loss of generality, assume $|a_{j}|\leq|b_{j}|$ and $|\sqrt{n_{j}+b_{j}}-\sqrt{n_{j}}|\leq c$, therefore
@@
$d_{j}$ $\coloneqq\sqrt{n_{j}+b_{j}}-\sqrt{n_{j}}-(\sqrt{n_{j}+a_{j}+b_{j}}-\sqrt{n_{j}+a_{j}})$
$\leq\sqrt{n_{j}+|b_{j}|}-\sqrt{n_{j}}-\Bigl{(}\sqrt{n_{j}+2|b_{j}|}-\sqrt{n_{j}+|b_{j}|}\Bigr{)}$
$\leq\sqrt{n_{j}+c\sqrt{n_{j}}}-\sqrt{n_{j}}-\Bigl{(}\sqrt{n_{j}+2c\sqrt{n_{j}}}-\sqrt{n_{j}+c\sqrt{n_{j}}}\Bigr{)}.$
@@
$d_{j}$ $\leq{n_{j}}^{1/4}\biggl{(}\frac{c}{2}\cdot{n_{j}}^{-1/4}+O\bigl{(}{n_{j}}^{-1/2}\bigr{)}\biggr{)}-{n_{j}}^{1/4}\biggl{(}\frac{c}{2}\cdot{n_{j}}^{-1/4}+O\bigl{(}{n_{j}}^{-1/2}\bigr{)}\biggr{)}$
$=\frac{c}{2}-\frac{c}{2}+O\bigl{(}{n_{j}}^{-1/4}\bigr{)}=o(1).$
We can now write
$u^{1}_{n_{j}}-u^{1}_{n_{j}+a_{j}}$ $=(\sqrt{n_{j}+b_{j}}-\sqrt{n_{j}})e^{2\pi i\alpha(n_{j}+b_{j})}$-
$\qquad(\sqrt{n_{j}+a_{j}+b_{j}}-\sqrt{n_{j}+a_{j}})e^{2\pi i\alpha(n_{j}+a_{j}+b_{j})}$
$=\bigl{(}(\sqrt{n_{j}+b_{j}}-\sqrt{n_{j}})(1-e^{2\pi i\alpha a_{j}})+d_{j}e^{2\pi i\alpha a_{j}}\bigr{)}e^{2\pi i\alpha(n_{j}+b_{j})},$
and since $\sqrt{n_{j}+b_{j}}-\sqrt{n_{j}}$ is bounded, $e^{2\pi i\alpha a_{j}}\to 1$, and $d_{j}\to 0$ we have that $u^{1}_{n_{j}}-u^{1}_{n_{j}+a_{j}}\to 0$.
@@
$\lim_{j\to\infty}u_{n_{j}+a_{j}}^{2}-u_{n_{j}}^{2}$ $=\lim_{j\to\infty}(x_{n_{j}+a_{j}}-x_{n_{j}})(e^{2\pi i\alpha b_{j}}-1)$
$=v_{1}\lim_{j\to\infty}(e^{2\pi i\alpha b_{j}}-1)=v_{1}(1-1)=0.$
To show that the Chabauty limit is closed under taking inverses, we can look at the points $x_{n_{j}-a_{j}}$, and if $x_{n_{j}-a_{j}}-x_{n_{j}}$ converges to $v$, then from the argument above we have that $v+v_{1}=0$, therefore $v_{1}=v$ is in the Chabauty limit. Note that we have
@@ 4 Properties of the Limit
[…]
@@
The spiral set is relatively dense (see [Aki20]), so there is $0<r$ such that for every $j$ there is $a_{j}$ such that $\|x_{n_{j}}-x_{n_{j}+a_{j}}\|<r$, therefore, by the definition of $q_{j}$, $\|x_{n_{j}}-x_{n_{j}+q_{j}}\|<r$, and the sequence $x_{n_{j}-q_{j}}-x_{n_{j}}$ must have a converging subsequence, say to a vector $v\neq 0$. Therefore, we have that a subsequence of $x_{n_{j}+q_{j}}-x_{n_{j}}$ converges to the vector $v$. Let $u\neq 0$ be some vector in the lattice, then there is a sequence $a_{j}$ such that $x_{n_{j}+a_{j}}-x_{n_{j}}\to u$, and by Proposition 8 for each $j$ we have
@@
Denote by $A$ the convex hull of $\pm v$ and $\pm\tilde{v}$. Assume that the lattice is not spanned by $v$ and $\tilde{v}$, so there is a vector $u\in\Lambda$ that is not spanned by them, and let $a_{j}$ be a sequence such that $x_{n_{j}+a_{j}}-x_{n_{j}}\to u$. Without loss of generality, we may assume that $u$ is in the basic parallelogram with the vertices at 0, $v$, $\tilde{v}$, and $v+\tilde{v}$. If $u\notin A$, then it is in the triangle with its vertices at $v$, $\tilde{v}$, and $v+\tilde{v}$, therefore $u-(v+\tilde{v})\in A$. This means that without loss of generality we may assume that $u\in A$ and we may also assume that $a_{j}>0$ because otherwise we can take $x_{n_{j}-a_{j}}-x_{n_{j}}\to-u\in A$.
@@
[…]
@@
[…]
@@
$\lim_{j\to\infty}v_{j}^{2}=\lim_{j\to\infty}e^{2\pi i\alpha n_{j}}\sqrt{n_{j}+q_{j}}(e^{2\pi i\alpha q_{j}}-1)=e^{i\theta}\cdot\lim_{j\to\infty}\sqrt{n_{j}+q_{j}}(e^{2\pi i\alpha q_{j}}-1).$
@@
$\mkern-36.0mu\operatorname{Re}\sqrt{n_{j}+q_{j}}(e^{2\pi i\alpha q_{j}}-1)=\sqrt{n_{j}+q_{j}}\cos(2\pi\alpha q_{j})$
$=\sqrt{n_{j}+q_{j}}O\biggl{(}\frac{1}{{q_{j}}^{2}}\biggr{)}=\sqrt{\frac{n_{j}}{{q_{j}^{2}}}+\frac{1}{q_{j}}}O\biggl{(}\frac{1}{q_{j}}\biggr{)}$
$=\biggl{(}\frac{n_{j}}{q_{j}}+o(1)\biggr{)}O\biggl{(}\frac{1}{q_{j}}\biggr{)}=(2t+o(1))o(1)=o(1)$
$\mkern-36.0mu\operatorname{Im}\sqrt{n_{j}+q_{j}}(e^{2\pi i\alpha q_{j}}-1)=\sqrt{n_{j}+q_{j}}\sin(2\pi\alpha q_{j})$
$=\sqrt{n_{j}+q_{j}}\biggl{(}2\pi(\alpha q_{j}-p_{j})+O\biggl{(}\frac{1}{{q_{j}}^{3}}\biggr{)}\biggr{)}$
$=\sqrt{\frac{n_{j}}{{q_{j}}^{2}}+\frac{1}{q_{j}}}\biggl{(}2\pi c+o(1)+O\biggl{(}\frac{1}{{q_{j}}^{2}}\biggr{)}\biggr{)}$
$=\biggl{(}\frac{\sqrt{n_{j}}}{q_{j}}+o(1)\biggr{)}\cdot(2\pi c+o(1))=\frac{1}{2t}\cdot 2\pi c+o(1)=\frac{\pi c}{t}+o(1),$
@@
[…]
@@
Proposition 12 clearly follows from the main result is given in [YSS21], where they prove that the area of Voronoi cells converges to $\pi$ (see the paper for definitions). However, we give a brief proof for completeness.
@@
$\frac{\tilde{c}}{\beta}-c\beta$ $=\bigl{(}\tilde{q}_{j}(\tilde{q}_{j}\alpha-\tilde{p}_{j})+o(1)\bigr{)}\biggl{(}\frac{q_{j}}{\tilde{q}_{j}}+o(1)\biggr{)}$-
$\qquad\bigl{(}q_{j}(q_{j}\alpha-p_{j})+o(1)\bigr{)}\biggl{(}\frac{\tilde{q}_{j}}{q_{j}}+o(1)\biggr{)}$
$=q_{j}(\tilde{q}_{j}\alpha-\tilde{p}_{j})-\tilde{q}_{j}(q_{j}\alpha-p_{j})+o(1),$
@@
$\biggl{|}\frac{\tilde{c}}{\beta}-c\beta\biggr{|}$ $=|q_{j}(\tilde{q}_{j}\alpha-\tilde{p}_{j})-\tilde{q}_{j}(q_{j}\alpha-p_{j})|+o(1)$
$=q_{j}|\tilde{q}_{j}\alpha-\tilde{p}_{j}|+\tilde{q}_{j}|q_{j}\alpha-p_{j}|+o(1)=1+o(1),$
@@
[…]
@@
$\mkern-36.0mu\operatorname{Re}e^{-2\pi i\alpha n_{j}}(x_{n_{j}+q_{j}}-x_{n_{j}})=\operatorname{Re}\sqrt{n_{j}+q_{j}}e^{2\pi i\alpha q_{j}}-\sqrt{n_{j}}$
$=\sqrt{n_{j}+q_{j}}\biggl{(}1-O\biggl{(}\frac{1}{{q_{j}}^{2}}\biggr{)}\biggr{)}-\sqrt{n_{j}}$
$=\sqrt{n_{j}+q_{j}}-\sqrt{n_{j}}+\sqrt{\frac{n_{j}}{{q_{j}}^{2}}+\frac{1}{q_{j}}}\cdot o(1)$
$=t+o(1)+\biggl{(}\frac{1}{2t}+o(1)\biggr{)}\cdot o(1)=t+o(1)$
$\mkern-36.0mu\operatorname{Im}e^{-2\pi i\alpha n_{j}}(x_{n_{j}+q_{j}}-x_{n_{j}})=\operatorname{Im}\sqrt{n_{j}+q_{j}}e^{2\pi i\alpha q_{j}}-\sqrt{n_{j}}$
$=\sqrt{n_{j}+q_{j}}\sin(2\pi\alpha q_{j})=\sqrt{n_{j}+q_{j}}\biggl{(}2\pi(q_{j}\alpha-p_{j})+O\biggl{(}\frac{1}{{q_{j}}^{3}}\biggr{)}\biggr{)}$
$=\sqrt{\frac{n_{j}}{{q_{j}}^{2}}+\frac{1}{q_{j}}}\biggl{(}2\pi c+O\biggl{(}\frac{1}{{q_{j}}^{2}}\biggr{)}\biggr{)}=\biggl{(}\frac{1}{2t}+o(1)\biggr{)}(2\pi c+o(1))=\frac{\pi c}{t}+o(1),$
@@
The author conjectures that there is a spiral set that is a dense forest, this is in contrast to a conjecture of Adiceam and Tsokanos in [AT22]. A possible avenue to consider is noting that if all of the Chabauty limits of a set are dense forests, the set is also a dense forest (by Proposition 6), and then constructing a spiral set whose limits are similar to the construction given by Bishop in [Bis11] using the lattice limits given by Fermat spirals. If there are indeed spiral sets that are dense forests, one might ask what is the best visibility function, and in particular whether there is a Danzer spiral set.